OpenStudy (anonymous):

-(7/5u) -(3/5) = 7u -(1/2) Simplified as much as possible. I think the answer is x = 85 Any help?

OpenStudy (texaschic101):

is your problem : -7/5(u) - 3/5 = 7u - 1/2 ?? OR -7/(5u) - 3/5 = 7u - 1/2

OpenStudy (anonymous):

let me use the equation option...one second

OpenStudy (anonymous):

\[-\frac{ 7 }{ 5 }u - \frac{ 3 }{ 5 } = 7u - \frac{ 1}{ 2 }\]

OpenStudy (anonymous):

It looks like that!

OpenStudy (texaschic101):

-7/5u - 3/5 = 7u - 1/2 --- multiply entire equation by the LCD, which is 10 -14u - 6 = 70u - 5 -- subtract 70u from both sides -14u - 70u - 6 = -5 -- add 6 to both sides -14u - 70u = -5 + 6 -- combine like terms -84u = 1 -- divide by -84 u = - 1/84 any questions ?

OpenStudy (anonymous):

ohhh ok, I see where I made the mistake

OpenStudy (anonymous):

thank you :)

OpenStudy (texaschic101):

anytime :)